Solve the system of equations using substitution for:
y = -5x + 22 and 2x = 0 +4y
Show solution to the system as a pair of coordinates.

Found 2 solutions by Alan3354, Collinwilfixyouuprealnice:
Answer by Alan3354(69443)

Solve the system of equations using substitution for:
y = -5x + 22
2x = 0 +4y --> x = 2y
-----
y = -5x + 22
y = -5*2y + 22
11y = 22
etc
Show solution to the system as a pair of coordinates.

Answer by Collinwilfixyouuprealnice(44)

2x-y=9;3x+4y=-14
Solution :
{x,y} = {2,-5}
Solve by Substitution :
Solve equation for the variable y

y = 2x - 9
Plug this in for variable y in equation
3x + 4•(2x-9) = -14
11x = 22
Solve equation for the variable x

11x = 22
x = 2
By now we know this much :
x = 2
y = 2x-9
Use the x value to solve for y
y = 2(2)-9 = -5
Solution :
{x,y} = {2,-5}